5 A particle is moving around in a circle of radius R = 1.5 m with a constant speed of 2 m/s. What is the centripetal acceleration and angular velocity of the particle?

5

A particle is moving around in a circle of radius R = 1.5 m with a constant speed of 2 m/s. What is the centripetal acceleration and angular velocity of the particle?

Grade:12

1 Answers

Nirmal Singh.
10 years ago
v = 2 m/s, R = 1.5 m,

centripetal acceleration, a = v2 / R = (2)2 / 1.5 = 4/1.5 = 2.65 m/s2

Angular velocity w = v/R = 2 /1.5 = 1.65 rad/s

Thanks & Regards,

Nirmal Singh

Askiitians Faculty